how much will you have in a savings account if you deposit $4000 for 8 years at 5% interest compounded annually?

Answer by mananth(16949)

Principal P = 4000
Amount= A
years=n 8.00
compounded 1 times a year t
Rate = 5.00 0.05
Amount = P*((n+r)/n)^n*t

Amount = = 4000 *( 1 + 0.05 )^ 8 * 1
Amount = 4000 *( 1 + 0.05 )^ 8
4000 *( 1.05 )^ 8
Amount = 5909.82
m.ananth@hotmail.ca